1.      Assume that in the decomposition of KCIO3 the volume of oxygen gas produced at STP in cm3 is 100 times its molar mass. Calculate the mass of KCI formed in this reaction 


1
Expert's answer
2022-05-12T17:51:03-0400

2KCIO= 2KCl + 3O2, n(O2)/3=n(KCl)/2=) n(KClO3)=n(O2)*1.5 =0.32*1.5/22.4 = 0.0214, m(KCl)=1.6 grams